11) A multiple choice test allows answers of a, b, c, d, e for each question. A student answers ten questions by making random guesses.

a) Find the probability that all ten answers are wrong. b) Find the probability that the first answer is correct.

No handwriting, please type your answer.
Answer #1

n=10, p=1/5=0.20, q=(1-p)=(1-0.20)=0.8

a) P(ten are wrong) = 10c0* (0.20)^0 *(0.80)^10

=0.1073

b) Here the question is about the first question alone hence n becomes 1

P(the first answer is correct) = p=0.20